# Necklaces, subset sums, and cyclic permutations

### 16 Mar 2026 v1

@@ Abstract
On the one hand, we introduce a parameter $r$ so that requiring the subset sums to be congruent to $r$ modulo $n$ translates into imposing some periodicity conditions on the necklaces. On the other hand, we refine these relations by the size $k$ of the subset, showing that it matches the number of ones in the necklace. We describe the precise conditions on $n$, $k$ and $r$ for which the equalities hold.
@@ 1 Introduction
As shown by Stanley [14, Exer. 1.105], the number of subsets of $n]\coloneqq\{1,2,\dots,n\}$ the sum of whose elements is divisible by $n$ is equal to
@@
One of such objects are primitive necklaces, which are necklaces all of whose rotations are different. It is known that equation (2) counts primitive necklaces of length $n$ with an odd number of ones. Another such object are V-shaped permutations of $[n]$ (i.e., permutations with a unique local minimum) which are cycles, also known [3] to be enumerated by equation (2). Baker et al. [4] noticed that, according to [12, A000048], there is yet another object counted by equation (2): subsets of [n1][n-1] whose elements sum to $1\bmod n$. They conjectured that the following refinement holds.
@@
Our first main result (Theorem 3.1 below) is a proof of this conjecture. We also show, by applying a bijection from [3, 16], that the connection that was observed in [4] between primitive necklaces and subsets that sum to $1\bmod n$ admits a natural refinement where the number of ones in the necklace corresponds to the size of the subset. The analogous refinement also applies to the equality from [14, Exer. 1.105] involving arbitrary necklaces and subsets whose sum is divisible by $n$.
@@
In Section 5, we list several open questions that our work raises, mostly about finding bijective proofs of several equalities relating necklaces, subset sums, and permutations.
@@ 3.1 A conjecture about subset sums and permutations
$\sum_{k=1}^{n}\left|S_{1}(n,k-1)\right|x^{k-1}=\frac{1}{n}\sum_{i=0}^{n-1}\zeta_{n}^{-i}f(x,\zeta_{n}^{i}).$ (6)
Fix $i$ and let $\omega=\zeta_{n}^{i}$, which is a root of unity of order $m\coloneqq n/{\gcd(n,i)}$. Write
$f(x,\zeta_{n}^{i})=f(x,\omega)=\prod_{j=1}^{n-1}(1+x\omega^{j}),$
@@
$f(x,\zeta_{n}^{i})=\frac{1}{1+x}\left(\prod_{s=0}^{m-1}(1+x\omega^{s})\right)^{n/m}.$
@@
$\zeta_{n}^{-i}f(x,\zeta_{n}^{i})=\frac{\zeta_{n}^{-i}}{1+x}\left(1-(-x)^{n/{\gcd(n,i)}}\right)^{\gcd(n,i)}.$
Plugging this back into equation (6) and grouping the terms of the sum where $\gcd(n,i)=d$, for each divisor $d$ of $n$, we obtain
$\sum_{k=1}^{n}\left|S_{1}(n,k-1)\right|x^{k-1}$ $=\frac{1}{n}\sum_{i=0}^{n-1}\frac{\zeta_{n}^{-i}}{1+x}\left(1-(-x)^{n/\gcd(n,i)}\right)^{\gcd(n,i)}$
$=\frac{1}{n(1+x)}\sum_{d\mid n}\left(1-(-x)^{n/d}\right)^{d}\left(\sum_{\begin{subarray}{c}i=0\\ \gcd(n,i)=d\end{subarray}}^{n-1}\zeta_{n}^{-i}\right).$
@@ 3.2 Bijections between permutations and necklaces
It follows from [16, Theorem 2.2] that $(w)$ is either primitive or of the form $(vv)$, where $v$ is primitive with an odd number, $k/2$, of ones. In the first case, we let $\Psi(\sigma)=(w)$; in the second case, which can only happen when $n$ is even and $k\equiv 2\bmod 4$, we let $\Psi(\sigma)=(v)$. In both cases, $\Psi(\sigma)\in L^{+}(n,k)$.
@@ 4.1 Formulas for arbitrary $r$
To compare the quantities $\left|\bar{S}_{r}(n,k)\right|$ and $\left|N_{r}(n,k)\right|$, we will first show that they can both be expressed in a similar form. These expressions involve the following sums of powers of a primitive $d$-root of unity $\zeta_{d}$, often called Ramanujan sums (see [2, Ch. 8.3], [1], or [12, A086831]):
@@
$\sum_{k=0}^{n}\left|\bar{S}_{r}(n,k)\right|x^{k}=\frac{1}{n}\sum_{i=0}^{n-1}\zeta_{n}^{-ri}\bar{f}(x,\zeta_{n}^{i}).$
Following the steps in that proof, and letting $m=n/\gcd(n,i)$, we can write
$\bar{f}(x,\zeta_{n}^{i})=\prod_{j=1}^{n}(1+x\zeta_{n}^{ij})=\prod_{s=0}^{m-1}\left(1+x\zeta_{n}^{is}\right)^{n/m}=\left(1-(-x)^{m}\right)^{n/m},$
@@
$\sum_{k=0}^{n}\left|\bar{S}_{r}(n,k)\right|x^{k}=\frac{1}{n}\sum_{d\mid n}\left(1-(-x)^{n/d}\right)^{d}\left(\sum_{\begin{subarray}{c}i=0\\ \gcd(n,i)=d\end{subarray}}^{n-1}\zeta_{n}^{-ri}\right).$
Because $\gcd(n,i)=d$, the root of unity $\zeta_{n}^{-i}$ has order $n/d$. Therefore the rightmost summation is the sum of the $r$th powers of roots of unity of order $n/d$, which is precisely $c_{n/d}(r)$. Replacing $d$ with $n/d$, we obtain
@@ 5 Open questions
Problem 5.3.
Find a bijection between $\bar{S}_{1}(n,k)$ and $L^{+}(n,k)$ for $n\geq 2$ and $0\leq k\leq n$. Interesting special cases include a bijection between $\bar{S}_{1}(n,k)$ and $L(n,k)$ any of the conditions (a)–(e) in Corollary 4.10 hold, or a bijection between $\bar{S}_{1}(n)$ and $L(n)$ when $n$ is odd.
Problems 5.2 and 5.3 can be generalized to arbitrary $r$ as follows.
Problem 5.4.
@@
Problem 5.5.
Are there analogues of the diagram in Figure 1 for arbitrary $r$? Specifially, are there sets of permutations playing the role of $\operatorname{CVP}(n,k)$?
In this paper, we have focused on binary necklaces and subsets, as they do in [4, 11, 13]. However, for any integer $q\geq 2$, one can define necklaces over the alphabet $\{0,1,\dots,q-1\}$, sometimes called $q$-ary or $q$-colored necklaces. In [6], Chan shows that, if $\gcd(q,n)=1$, then the number of $q$-ary necklaces of length $n$ equals the number of multisets of $[n]$ whose sum is divisible by $n$ and where the multiplicity of each element is less than $q$. This generalizes the fact that $\left|\bar{S}_{0}(n)\right|=\left|N(n)\right|$ for odd $n$, as discussed in Section 2.4.
It is natural to ask whether our generalization involving the parameter $r$, as well as the refinement by the parameter $k$, can be extended from binary necklaces to $q$-ary necklaces. More specifically, for any integers $0\leq k\leq n$, $0\leq r<n$, and $q\geq 2$, one can define the following sets:
Computational evidence suggests that these sets often have the same cardinality.
Problem 5.6.
Generalize our results to $q$-ary necklaces. Specifically, describe necessary and sufficient conditions on $n,k,r,q$ for which the equality $\left|\bar{S}_{r}^{q}(n,k)\right|=\left|N_{r}^{q}(n,k)\right|$ holds.
Theorem 4.7 describes these conditions when $q=2$. On the other hand, in the case $r=0$ and without the refinement by $k$, Chan [6] proves that, if $\gcd(q,n)=1$, then $\left|\bar{S}_{0}^{q}(n)\right|=\left|N_{0}^{q}(n)\right|$ (defined as in equation (3)), and gives a bijective proof of this equality when $q$ is a prime power. Computational data suggests that the equality $\left|\bar{S}_{0}^{q}(n,k)\right|=\left|N_{0}^{q}(n,k)\right|$ may hold for all $k$, provided that $\gcd(q,n)=1$.

### 21 Apr 2026 v2

@@ Abstract
On the one hand, we introduce a parameter $r$ so that requiring the subset sums to be congruent to $r$ modulo $n$ translates into imposing some periodicity conditions on the necklaces. On the other hand, we refine these relations by the size $k$ of the subset, showing that it matches the number of ones in the necklace. We describe the precise conditions on $n$, $k$ and $r$ for which the equalities hold. We also extend some of our formulas to $q$-ary necklaces.
@@ 1 Introduction
As shown by Stanley [15, Exer. 1.105], the number of subsets of $n]\coloneqq\{1,2,\dots,n\}$ the sum of whose elements is divisible by $n$ is equal to
@@
One of such objects are primitive necklaces, which are necklaces all of whose rotations are different. It is known that equation (2) counts primitive necklaces of length $n$ with an odd number of ones. Another such object are V-shaped permutations of $[n]$ (i.e., permutations with a unique local minimum) which are cycles, also known [3] to be enumerated by equation (2). Baker et al. [4] noticed that, according to [13, A000048], there is yet another object counted by equation (2): subsets of [n1][n-1] whose elements sum to $1\bmod n$. They conjectured that the following refinement holds.
@@
Our first main result (Theorem 3.1) is a proof of this conjecture. We also show, by applying a bijection from [3, 17], that the connection that was observed in [4] between primitive necklaces and subsets that sum to $1\bmod n$ admits a natural refinement where the number of ones in the necklace corresponds to the size of the subset. The analogous refinement also applies to the equality from [15, Exer. 1.105] involving arbitrary necklaces and subsets whose sum is divisible by $n$.
@@
In Section 5 we extend some of our results to $q$-ary necklaces and multisets, refining and generalizing an identity found by Chan [6]. Finally, in Section 6, we list several open questions that our work raises, mostly about finding bijective proofs of several equalities relating necklaces, subset sums, and permutations.
@@ 3.1 A conjecture about subset sums and permutations
$\sum_{k=1}^{n}\left|S_{1}(n,k-1)\right|x^{k-1}=\frac{1}{n}\sum_{\ell=0}^{n-1}\zeta_{n}^{-\ell}f(x,\zeta_{n}^{\ell}).$ (6)
Fix $\ell$ and let $\omega\coloneqq\zeta_{n}^{\ell}$, which is a root of unity of order $m\coloneqq n/{\gcd(n,\ell)}$. Write
$f(x,\zeta_{n}^{\ell})=f(x,\omega)=\prod_{j=1}^{n-1}(1+x\omega^{j}),$
@@
$f(x,\zeta_{n}^{\ell})=\frac{1}{1+x}\left(\prod_{s=0}^{m-1}(1+x\omega^{s})\right)^{n/m}.$
@@
$\zeta_{n}^{-\ell}f(x,\zeta_{n}^{\ell})=\frac{\zeta_{n}^{-\ell}}{1+x}\left(1-(-x)^{n/{\gcd(n,\ell)}}\right)^{\gcd(n,\ell)}.$
Plugging this back into equation (6) and grouping the terms of the sum where $\gcd(n,\ell)=d$, for each divisor $d$ of $n$, we obtain
$\sum_{k=1}^{n}\left|S_{1}(n,k-1)\right|x^{k-1}$ $=\frac{1}{n}\sum_{\ell=0}^{n-1}\frac{\zeta_{n}^{-\ell}}{1+x}\left(1-(-x)^{n/\gcd(n,\ell)}\right)^{\gcd(n,\ell)}$
$=\frac{1}{n(1+x)}\sum_{d\mid n}\left(1-(-x)^{n/d}\right)^{d}\left(\sum_{\begin{subarray}{c}\ell=0\\ \gcd(n,\ell)=d\end{subarray}}^{n-1}\zeta_{n}^{-\ell}\right).$
@@ 3.2 Bijections between permutations and necklaces
It follows from [17, Thm. 2.2] that $(w)$ is either primitive or of the form $(vv)$, where $v$ is primitive with an odd number, $k/2$, of ones. In the first case, we let $\Psi(\sigma)=(w)$; in the second case, which can only happen when $n$ is even and $k\equiv 2\bmod 4$, we let $\Psi(\sigma)=(v)$. In both cases, $\Psi(\sigma)\in L^{+}(n,k)$.
@@ 4.1 Formulas for arbitrary $r$
To compare the quantities $\left|\bar{S}_{r}(n,k)\right|$ and $\left|N_{r}(n,k)\right|$, we will first show that they can both be expressed in a similar form. These expressions involve the following sums of powers of a primitive $d$-root of unity $\zeta_{d}$, often called Ramanujan sums (see [2, Ch. 8.3], [1], or [13, A086831]):
@@
$\sum_{k=0}^{n}\left|\bar{S}_{r}(n,k)\right|x^{k}=\frac{1}{n}\sum_{\ell=0}^{n-1}\zeta_{n}^{-r\ell}\bar{f}(x,\zeta_{n}^{\ell}).$
Following the steps in that proof, and letting $m=n/\gcd(n,\ell)$, we can write
$\bar{f}(x,\zeta_{n}^{\ell})=\prod_{j=1}^{n}(1+x\zeta_{n}^{\ell j})=\left(\prod_{s=0}^{m-1}(1+x\zeta_{n}^{\ell s})\right)^{n/m}=\left(1-(-x)^{m}\right)^{n/m},$
@@
$\sum_{k=0}^{n}\left|\bar{S}_{r}(n,k)\right|x^{k}=\frac{1}{n}\sum_{d\mid n}\left(1-(-x)^{n/d}\right)^{d}\left(\sum_{\begin{subarray}{c}\ell=0\\ \gcd(n,\ell)=d\end{subarray}}^{n-1}\zeta_{n}^{-r\ell}\right).$
Because $\gcd(n,\ell)=d$, the root of unity $\zeta_{n}^{-\ell}$ has order $n/d$. Therefore the rightmost summation is the sum of the $r$th powers of roots of unity of order $n/d$, which is precisely $c_{n/d}(r)$. Replacing $d$ with $n/d$, we obtain
@@ 4.4 Special cases
5 Generalizations to $q$-ary necklaces
This paper has focused on binary necklaces and subsets, following the approach from [4, 12, 14]. However, for any integer $q\geq 2$, one can define necklaces over the alphabet $\{0,1,\dots,q-1\}$, sometimes called $q$-ary or $q$-colored necklaces. In [6], Chan shows that, if $\gcd(q,n)=1$, then the number of $q$-ary necklaces of length $n$ equals the number of multisets of $[n]$ whose sum is divisible by $n$ and where the multiplicity of each element is less than $q$. This generalizes the fact that $\left|\bar{S}_{0}(n)\right|=\left|N(n)\right|$ for odd $n$, as discussed in Section 2.4. In fact, Chan gives a bijective proof of his equality when $q$ is a prime power. This bijection has recently been generalized by Li and Zhou [10] to any $q$ such that $\gcd(q,n)=1$.
In this section we briefly discuss how some of our generalized results involving the parameters $r$ and $k$ can be extended from binary necklaces to $q$-ary necklaces. For this extension to work, the interpretation of the parameter $k$ that we use on $q$-ary necklaces is the sum of the entries, viewed as integers between 0 and q1q-1.
Definition 4.
For any integers $0\leq k\leq n$, $0\leq r<n$, and $q\geq 2$, define the following sets:
Next we generalize Lemmas 4.1 and 4.2 to arbitrary $q$. It is more convenient to give expressions for the generating functions where we sum over $k$.
Theorem 5.1.
For any $0\leq r<n$ and any $q\geq 2$,
$\sum_{k=0}^{n}\left|\bar{S}_{r}^{q}(n,k)\right|x^{k}=\frac{1}{n}\sum_{d\mid n}c_{d}(r)\left(\frac{\left(1-x^{\frac{qd}{\gcd(q,d)}}\right)^{\gcd(q,d)}}{1-x^{d}}\right)^{n/d}.$
Proof.
Generalizing the proof of Lemma 4.1, consider now the generating function
$\bar{f}_{q}(x,t)=\prod_{j=1}^{n}(1+xt^{j}+x^{2}t^{2j}+\dots+x^{q-1}t^{(q-1)j})=\prod_{j=1}^{n}\frac{1-x^{q}t^{qj}}{1-zt^{j}}$
for multisets of $[n]$ where each element has multiplicity less than $q$, with respect to their size and their sum. Letting $\zeta_{n}$ be a primitive $n$th root of unity, we have
$\sum_{k=0}^{n}\left|\bar{S}_{r}^{q}(n,k)\right|x^{k}=\frac{1}{n}\sum_{\ell=0}^{n-1}\zeta_{n}^{-r\ell}\bar{f}_{q}(x,\zeta_{n}^{\ell}).$ (17)
For any given $\ell$, we can evaluate $\bar{f}_{q}(x,\zeta_{n}^{\ell})$ directly:
$\bar{f}_{q}(x,\zeta_{n}^{\ell})=\prod_{j=1}^{n}(1-x^{q}\zeta_{n}^{\ell qj})\prod_{j=1}^{n}(1-z\zeta_{n}^{\ell j})^{-1}.$
For the first product, if we set $m\coloneqq n/\gcd(n,\ell q)$, then $w\coloneqq\zeta_{n}^{\ell q}$ is a root of unity of order $m$, so the product is
$\prod_{j=1}^{n}(1-x^{q}w^{j})=\left(\prod_{s=0}^{m-1}(1-x^{q}w^{s})\right)^{n/m}=\left(1-x^{qm}\right)^{n/m},$
as in the proof of Lemma 4.1.
The second product is handled similarly, with $m^{\prime}\coloneqq n/\gcd(n,\ell)$ and $w^{\prime}\coloneqq\zeta_{n}^{\ell}$ playing the roles of $m$ and $w$. We obtain
$\bar{f}_{q}(x,\zeta_{n}^{\ell})=\frac{\left(1-x^{qm}\right)^{n/m}}{\left(1-x^{m^{\prime}}\right)^{n/m^{\prime}}}=\frac{\left(1-x^{\frac{qn}{\gcd(n,\ell q)}}\right)^{\gcd(n,\ell q)}}{\left(1-x^{\frac{n}{\gcd(n,\ell)}}\right)^{\gcd(n,\ell)}}.$
Letting $d=\gcd(n,\ell)$, and noting that $\gcd(n,\ell q)=d\,\gcd\left(n/d,q\right)$, we can write
$\bar{f}_{q}(x,\zeta_{n}^{\ell})=\left(\frac{\left(1-x^{\frac{qn/d}{\gcd(n/d,q)}}\right)^{\gcd(n/d,q)}}{1-x^{n/d}}\right)^{d}.$
Plugging this into equation (17) and grouping according to $d=\gcd(n,\ell)$, we obtain
$\sum_{k=0}^{n}\left|\bar{S}_{r}^{q}(n,k)\right|x^{k}=\frac{1}{n}\sum_{d\mid n}\left(\frac{\left(1-x^{\frac{qn/d}{\gcd(n/d,q)}}\right)^{\gcd(n/d,q)}}{1-x^{n/d}}\right)^{d}\left(\sum_{\begin{subarray}{c}\ell=0\\ \gcd(n,\ell)=d\end{subarray}}^{n-1}\zeta_{n}^{-r\ell}\right).$
As before, the inner sum equals $c_{n/d}(r)$. Finally, if we replace $d$ with $n/d$, we obtain the stated expression. ∎
Theorem 5.2.
For any $0\leq r<n$ and any $q\geq 2$,
$\sum_{k=0}^{n}\left|N_{r}^{q}(n,k)\right|x^{k}=\frac{1}{n}\sum_{d\mid n}c_{d}(r)\left(\frac{1-x^{qd}}{1-x^{d}}\right)^{n/d}$
Proof.
For any $q$-tuple of nonnegative integers $\mathbf{n}=(n_{0},n_{1},\dots,n_{q-1})$, let $M_{r}(\mathbf{n})$ be the number of $q$-ary necklaces of length $|\mathbf{n}|=n_{0}+n_{1}+\cdots+n_{q-1}$ with $n_{i}$ copies of $i$ for every $i$, and whose co-period divides $r$. This is related to our generating function via the equation
$\sum_{k=0}^{n}\left|N_{r}^{q}(n,k)\right|x^{k}=\sum_{|\mathbf{n}|=n}M_{r}(\mathbf{n})\,x^{n_{1}+2n_{2}+\cdots+(q-1)n_{q-1}}.$ (18)
In [9, Thm. 1, eq. (18)], Hadjicostas shows that
$\sum_{\mathbf{n}\neq\mathbf{0}}M_{r}(\mathbf{n})\prod_{i=0}^{q-1}x_{i}^{n_{i}}=-\sum_{d\geq 1}\frac{c_{d}(r)}{d}\log(1-x_{0}^{d}-x_{1}^{d}-\cdots-x_{q-1}^{d}).$
Setting $x_{i}=yx^{i}$ for all $i$, this equation becomes
$\sum_{\mathbf{n}\neq\mathbf{0}}M_{r}(\mathbf{n})\,y^{|\mathbf{n}|}x^{n_{1}+2n_{2}+\cdots+(q-1)n_{q-1}}=-\sum_{d\geq 1}\frac{c_{d}(r)}{d}\log\left(1-y^{d}\frac{1-x^{qd}}{1-x^{d}}\right).$
Extracting the coefficient of $y^{n}$ on both sides, using equation (18) and the series expansion $\log(1-z)=\sum_{j\geq 1}z^{j}/j$, we obtain
$\sum_{k=0}^{n}\left|N_{r}^{q}(n,k)\right|x^{k}=\sum_{d\mid n}\frac{c_{d}(r)}{d}\frac{1}{n/d}\left(\frac{1-x^{qd}}{1-x^{d}}\right)^{n/d}=\frac{1}{n}\sum_{d\mid n}c_{d}(r)\left(\frac{1-x^{qd}}{1-x^{d}}\right)^{n/d}.\qed$
The following result generalizes Theorem 4.6 in the case of odd $n$.
Corollary 5.3.
If $\gcd(q,n)=1$, then
$\left|\bar{S}_{r}^{q}(n,k)\right|=\left|N_{r}^{q}(n,k)\right|.$
Proof.
If $\gcd(q,n)=1$, then $\gcd(q,d)=1$ for all $d\mid n$, and the formulas in Theorems 5.1 and 5.2 are identical in this case. ∎
@@ 6 Open questions
Problem 6.3.
Find a bijection between $\bar{S}_{1}(n,k)$ and $L^{+}(n,k)$ for $n\geq 2$ and $0\leq k\leq n$. Interesting special cases include a bijection between $\bar{S}_{1}(n,k)$ and $L(n,k)$ when any of the conditions (a)–(e) in Corollary 4.10 hold, or a bijection between $\bar{S}_{1}(n)$ and $L(n)$ when $n$ is odd.
Problems 6.2 and 6.3 can be generalized to arbitrary $r$ as follows.
Problem 6.4.
@@
Problem 6.5.
Are there analogues of the diagram in Figure 1 for arbitrary $r$? Specifically, are there sets of permutations playing the role of $\operatorname{CVP}(n,k)$?
Our last problem concerns the generalization of Theorem 4.7 to $q$-ary necklaces.
Problem 6.6.
Describe necessary and sufficient conditions on $n,k,r,q$ for which the equality $\left|\bar{S}_{r}^{q}(n,k)\right|=\left|N_{r}^{q}(n,k)\right|$ holds.
Corollary 5.3 gives the sufficient condition $\gcd(q,n)=1$.